# Blueberry galaxies up to 200 Mpc and their optical/infrared properties

### v1

@@ 5 Conclusions
In order to identify Blueberry galaxies and study their properties probed by infrared and visible light, this study utilized the HECATE catalog, the Pan-STARRS and SDSS photometries, the MPA-JHU spectroscopic analysis, and the ALLWISE photometry. In summary this work:
has identified 48 BBs, 40 of which are newly discovered. Most are closer compared to the previously known and 14 are in the south equatorial sky;
has reported the distribution of the BB’s physical size and showed that BBs have lower metallicities and stronger extinction compared to the expected when accounting for their and SFR;
has shown that BBs are at the extreme end of dwarf SFGs. Compared to GPs, BBs are on average less massive, and form stars more intensively, showing more strongly ionized ISM, lower metallicities, intrinsically bluer visible colors, and redder IR colors;
showed that BDGs with higher specific SFR tend to have redder IR colors with the correlation being stronger with the color.
@@ Appendix A The selection algorithm
The following Equation LABEL:eq:Selection shows the algorithm used for the selection of Blueberry galaxies, as described in Section 2.
Appendix B The SED fitting process
Figure 12 shows two representative SED fits performed with CIGALE, the worst (BB 20) and one of the best (BB 30) based on . Figures 13, and 14 show the -weighted probability densities, accounting all fitted models, of the , and SFR respectively for the 48 BBs of this work. The SFR and quantities were estimated based on the modes and 68% C.Is. of these distributions.

### v3

@@ 5 Conclusions
In order to identify BB galaxies and study their properties probed by infrared and visible light, this study utilized the HECATE catalog, the Pan-STARRS and SDSS photometries, the MPA-JHU spectroscopic analysis, and ALLWISE photometry. In summary, this work has done the following:
Identified 48 BBs, 40 of which are first recognized as such from this work. Most are closer compared to the previously known BBs, and 14 are in the south equatorial sky.
Reported the distribution of the BBs’ physical sizes and showed that BBs have lower metallicities and stronger extinction compared to the expected values when accounting for their and SFRs.
Showed that BBs are at the extreme end of dwarf SFGs. Compared to GPs, BBs are on average less massive, and they form stars more intensively, showing more strongly ionized ISM, lower metallicities, intrinsically bluer visible colors, and redder IR colors.
Showed that BDGs with higher specific SFRs tend to have redder IR colors, with the correlation being stronger with the color.
This sample of BBs can be used as a basis for future multi-wavelength studies of some of the most highly star-forming dwarf galaxies of the local Universe. Moreover, since BBs and GPs are the best local analogs of the dwarf SFGs of the early Universe, the results of this work provide constraints regarding their optical and IR emission.
@@ Acknowledgements.
- Cutri et al. (2021) Cutri, R. M., Wright, E. L., Conrow, T., et al. 2021, VizieR Online Data Catalog, II/328
@@
- Lumbreras-Calle et al. (2022) Lumbreras-Calle, A., López-Sanjuan, C., Sobral, D., et al. 2022, A&A, 668, A60
@@ Appendix A The selection algorithm
The following Equation LABEL:eq:Selection shows the algorithm used for the selection of BB galaxies (as described in Section 2):
Appendix B The spectral energy distribution fitting process
Figure 12 shows two representative SED fits performed with CIGALE, namely the worst (BB 20) and one of the best (BB 30), based on . Figures 13 and 14 show the -weighted probability densities, accounting for all fitted models, of the and SFR, respectively, for the 48 BBs of this work. The SFR and quantities were estimated based on the modes and 68% CIs of these distributions.